If there is a 20% chance we will get a 16% return, a 30% chance of getting a 14% return, a 40% chance of getting a 12% return, and a 10% chance of getting an 8% return.

Question 1: What is the expected rate of return for this investment?  (Please round your answer to the fourth decimal. E.g. if your answer is 1.11%, you should input 0.0111.)

Question 2: What is the standard deviation for this investment?  (Please round your answer to the fourth decimal. E.g. if your answer is 1.11%, you should input 0.0111.)
Answer #1

Answer 1.

Expected Return = 0.20 * 0.16 + 0.30 * 0.14 + 0.40 * 0.12 + 0.10 * 0.08
Expected Return = 0.13 or 13.00%

Answer 2.

Variance = 0.20 * (0.16 - 0.13)^2 + 0.30 * (0.14 - 0.13)^2 + 0.40 * (0.12 - 0.13)^2 + 0.10 * (0.08 - 0.13)^2
Variance = 0.0005

Standard Deviation = (0.0005)^(1/2)
Standard Deviation = 0.0224 or 2.24%